5.75 g of LiCI is dissolved in enough water 0.970
Answer #1

1) moles of LiCl in 5.75 g of it = mass/molar mass = 5.75/42.394 = 0.136

Thus, molarity of LiCl solution = moles/volume of solution in litres = 0.136/0.97 = 0.14 M

2) Number of Moles = molarity*volume of solution in litres = 0.14*0.01 = 0.0014

3) volume of solution = moles/molarity = 0.0078/0.14 = 0.0557 litres = 55.7 ml

4) CH3CH2OH can form H-bonds